OpenStudy (anonymous):

The equation of a least-squares regression line is 10+5x . If x increases by one unit, what is the corresponding increase in y ? A. 10 B. 5 C. 1 D. None of the above.

OpenStudy (anonymous):



OpenStudy (anonymous):

use plugging in values :D

OpenStudy (anonymous):

how do i plug in the values?

OpenStudy (anonymous):

f(x)=10+5x f(x+1)=10+5(x+1)=15+5x f(x+1)-f(x)=5

OpenStudy (auctoratrox):

lets say we use x = 1 and x = 2, because the question calls for x to increase by 1 unit, so now lets substitute the x-values into y = 10+5x, we get 15 and then 20, therefore we can say that for our particular regression line, for every change in x for 1 unit, the change in y is 5 units, which is B.